what is the future value if a 1000 annuity payment over 4 years if thrninterest rates are 18%
Answer #1

Future Value Annuity = (1+r)n-1 Periodic Payment *

where r is the rate of Return for compounding period = 18% OR 0.18

n is the no of compounding period 4 years

= Tomm (1+0.1814 -1 018

= 1000 * 5.215432

= $ 5215.432  

Future Value of 1000 Annuity will be $ 5215.432

NOTE: The answer to your question has been given below/above. If there is any query regarding the answer, please ask in the comment section. If you find the answer helpful, do upvote. Help us help you.